# A Unifying Framework for Parallelizing Sequential Models with Linear Dynamical Systems

### v1

@@ Abstract
Harnessing parallelism in seemingly sequential models is a central challenge for modern machine learning. Several approaches have been proposed for evaluating sequential processes in parallel using fixed-point methods, like Newton, Picard, and Jacobi iterations. In this work, we show that these methods can be understood within a common framework based on linear dynamical systems (LDSs), where different iteration schemes arise naturally as approximate linearizations of a nonlinear recursion. This unifying view highlights shared principles behind these techniques and clarifies when particular fixed-point methods are most likely to be effective. By bridging diverse algorithms through the language of LDSs, our framework provides a clearer theoretical foundation for parallelizing sequential models and points toward new opportunities for efficient and scalable computation.
@@ 1 Introduction
Sequential processes are ubiquitous in machine learning models. Evaluating a recurrent neural network (Goodfellow et al., 2016), sampling a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), generating from a deep state space model (Gu et al., 2022; Smith et al., 2023; Orvieto et al., 2023; Gu & Dao, 2024), and unrolling layers of a deep neural network (He et al., 2016; Vaswani et al., 2017) all involve sequential computations. Naively, these computations require time proportional to the length of the input or the depth of the architecture, and in some cases, they may not take full advantage of hardware accelerators like GPUs and TPUs (Lim et al., 2024). However, certain sequential computations — namely, linear recursions or linear dynamical systems (LDSs) — can be evaluated in parallel using techniques like the parallel scan (Blelloch, 1990; Fatahalian & Olukotun, 2024). Indeed, the parallelizability of linear recursions is key to efficiently evaluating many deep state space models (Smith et al., 2023; Gu & Dao, 2024). A natural question is whether other sequential processes in machine learning, which typically correspond to nonlinear recursions, can be similarly accelerated.
On first inspection, the parallel scan algorithm does not seem to generalize to nonlinear recursions. The reason it applies to linear recursions is that the composition of two linear functions remains linear, whereas the composition of two nonlinear functions is generally more complicated. For example, the composition of two quadratic functions is a quartic function. Nevertheless, recent works have proposed several techniques to parallelize nonlinear recursions, including Jacobi (Song et al., 2021a), Picard (Shih et al., 2023), and Newton (Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024) iterations. These techniques were originally applied to different machine learning problems and follow different notations and intuitions, which obscures their underlying similarities.
In this paper, we show that Jacobi, Picard, and Newton iterations all solve a fixed-point problem by iteratively linearizing the nonlinear recurrence and evaluating the resulting linear dynamical system with a parallel scan. In addition to providing a unifying perspective on these methods, we discuss the properties of each approach and their applicability for different types of problems. While the general connections between Picard and Newton iterations and their convergence rates for solving nonlinear equations have long been known by the applied mathematics community (Ortega & Rheinboldt, 1970), the novel contribution of this work is emphasizing the tight connection between these approaches for the specific problem of evaluating a nonlinear recursion, a broadly important problem in machine learning. In particular, we show how each fixed-point iteration reduces to evaluating an LDS, which crucially allows the use of parallel computation in what at first seems to be an inherently sequential problem.
@@
  • In Section˜2, we show how Newton, quasi-Newton, Picard, and Jacobi iterations for evaluating nonlinear recursions can all be cast under a unifying framework of iterative evaluations of LDSs.
  • In Section˜3, we discuss how many important problems in machine learning can be framed as evaluating a nonlinear recursion, and we illustrate how the structure of those problems informs which fixed-point methods are best suited for the task.
These contributions unify and clarify several recently proposed methods in the machine learning literature, and they highlight the central importance of linear dynamical systems for parallelizing seemingly sequential processes.
@@ 2 Unifying fixed-point iterations using linear dynamical systems
We first introduce notation for evaluating a generic sequence. Let $x_{t}\in\mathbb{R}^{D}$ denote the state at time $t$, and let $f_{t}$ denote the corresponding transition function at that time point. Throughout this paper, we will use $D$ to denote the dimension of the hidden state and $T$ to denote the sequence length.
@@
In order to be a fixed-point operator, $\mathcal{A}$ should have a unique fixed point $\mathbf{x}_{1:T}^{\star}$, which is the sequential roll-out of eq.˜1. Equivalently, $\mathcal{A}$ should satisfy that $\mathbf{x}_{1:T}^{\star}=\mathcal{A}\left(\mathbf{x}_{1:T}^{\star}\right)$, where $\mathbf{x}_{1:T}^{\star}$ is the unique root of the system of equations given by
$x_{t+1}-f_{t+1}(x_{t})=0\quad\forall t\in\{0,\ldots,T-1\}.$ (3)
@@
$x_{t+1}^{(i+1)}=f_{t+1}(x_{t}^{(i)})+A_{t+1}\,(x_{t}^{(i+1)}-x_{t}^{(i)}),$ (4)
where the transition matrix $A_{t+1}\in\mathbb{R}^{D\times D}$ is determined by the dynamics functions $f_{t+1}$ and the current guess for the state $x_{t}^{(i)}$. Different fixed-point methods use different transition matrices, as shown in Table˜1.
The transition matrix $A_{t+1}$ can be thought of as an approximation to the Jacobian of the dynamics function $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. Different fixed-point methods simply use different approaches to linearizing the dynamics $f_{t+1}$. Importantly, because the recursion in eq.˜4 is an LDS, it can be evaluated with a parallel scan, which runs in $\mathcal{O}(\log T)$ time on a machine with $\mathcal{O}(T)$ processors (Blelloch, 1990). We provide an introduction to the parallel scan algorithm in Appendix˜B.
@@
Table 1: Summary of fixed-point iteration schemes as linear dynamical systems. We order the methods by the number of fixed-point iterations in which they converge (fewer iterations needed to more iterations needed). However, it is important to note that the increased rate of convergence is offset by increased computational and memory requirements of each method. Each iteration is an LDS, i.e. can be written in the form of eq.˜4, where $A_{t+1}$ is the transition matrix of the LDS, and an approximation to the Jacobian of the dynamics function $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. See Algorithm˜1 as well. Throughout we use subscript $t$ to denote sequence position and superscript $(i)$ to indicate fixed-point iteration number. All of these methods are guaranteed to converge in at most $T$ iterations (Gonzalez et al., 2024, Proposition 1). We use “order” to mean the highest number of derivatives taken: Newton and quasi-Newton methods use first derivatives, while Picard and Jacobi methods do not use derivatives of $f_{t}$.
@@
Fixed-point method Order Transition matrix $A_{t+1}$
Newton first-order $\frac{\partial f_{t+1}}{\partial x_{t}}\!\left(x_{t}^{(i)}\right)$
@@
Quasi-Newton quasi first-order $\operatorname{diag}\left[\frac{\partial f_{t+1}}{\partial x_{t}}\left(x_{t}^{(i)}\right)\right]$
Picard zeroth-order $I_{D}$
Jacobi zeroth-order $0$
@@
In the rest of this section, we discuss in more detail how the four prominent fixed-point methods discussed in Table˜1 reduce to iterative application of LDSs when used to solve a recursion. Fundamentally, all of these methods parallelize nonlinear recursions by iteratively linearizing and evaluating them, as we indicate in Algorithm˜1. However, the different methods use different approximations of the Jacobian $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ of the dynamics function for their transition matrices $A_{t+1}$.
@@
   $A_{1:T}\leftarrow\textsc{LinearizeDynamics}(f,x_{0},\mathbf{x}_{1:T}^{(i)})$ $\triangleright$ For all $t$ in parallel
   $\mathbf{x}_{1:T}^{(i+1)}\leftarrow\textsc{EvaluateLDS}(x_{0},A_{1:T},f,\mathbf{x}_{1:T}^{(i)})$ $\triangleright$ Using parallel scan
@@ 2.1 Newton iterations
Danieli et al. (2023), Lim et al. (2024) and Gonzalez et al. (2024) demonstrated that when the fixed-point operator $\mathcal{A}$ is constructed as an appropriately designed LDS, built from a linearization of original nonlinear recursion $f$, then each application of $\mathcal{A}$ is equivalent to an iteration of Newton’s root-finding method on the system of equations given in eq.˜3. Specifically, each Newton fixed-point iteration, $\mathbf{x}_{1:T}^{(i+1)}=\mathcal{A}_{\mathrm{N}}\big(\mathbf{x}_{1:T}^{(i)}\big)$}, is defined by the linear recursion,
@@
which we recognize as first-order Taylor expansion of the nonlinear recursion. Because this fixed-point iteration uses a first derivative, we refer to it as a first-order fixed-point iteration. We note that for Newton iterations, the transition matrix is exactly the Jacobian of the dynamics function, $A_{t+1}=\tfrac{\partial f_{t+1}}{\partial x_{t}}(x_{t}^{(i)})$.
Because eq.˜5 is a linear dynamical system, it can be evaluated with a parallel scan. However, each iteration requires $\mathcal{O}(TD^{2})$ memory to store the $T$ Jacobian matrices, and $\mathcal{O}(TD^{3})$ work to compute the matrix-matrix multiplies in the parallel scan. This expense is prohibitive for large state size or sequence length, which motivates the quasi-Newton iterations we discuss next.
@@ 2.2 Quasi-Newton iterations
Because full Newton iterations are costly in compute and memory, there is a wide literature on quasi-Newton methods (Nocedal & Wright, 2006). A particularly simple way to turn the Newton iteration in eq.˜5 into a quasi-Newton iteration that uses a parallel associative scan was proposed by Gonzalez et al. (2024): just use the diagonal of the Jacobian of the dynamics function. Specifically, each of these quasi-Newton fixed-point iterations $\mathcal{A}_{\mathrm{QN}}$ is given by
@@
which we recognize as an LDS with transition matrix $A_{t+1}=\mathrm{diag}\big[\tfrac{\partial f_{t+1}}{\partial x}(x_{t}^{(i)})\big]$.
With this transition matrix, the parallel scan requires only $\mathcal{O}(TD)$ space and $\mathcal{O}(TD)$ work, and is therefore more computationally efficient than the full Newton iteration eq.˜5. However, quasi-Newton methods take more fixed-point iterations to converge. As discussed by Gonzalez et al. (2024) and Zoltowski et al. (2025) and shown in our experiments in Section˜3, whether Newton and quasi-Newton methods run faster for evaluating a given sequence is an empirical question whose answer depends on many factors, including choice of hardware, scale of the problem, and the faithfulness of the approximate Jacobian to the true dynamics. However, the large memory consumption of full Newton iterations renders it infeasible for large-scale problems.
In general, any structured approximation of the Jacobian matrix that remains closed under matrix multiplication [^1] Closed in the sense that the product, $A_{t+1}A_{t}$, has the same structure as $A_{t+1}$ and $A_{t}$, as with diagonal matrices. See Appendix B for more detail. could be used to form a parallel quasi-Newton fixed-point iteration for evaluating a sequence. For example, Zoltowski et al. (2025) introduces a parallel quasi-Newton method where each block of the matrix is diagonal. The broader development of quasi-Newton methods that fit into the unifying LDS framework discussed in this paper is an important direction for future work, which we elaborate on in Section˜5. However, for simplicity, henceforth in this paper when we refer to “quasi-Newton iterations,” we restrict ourselves to the simple diagonal approximation shown in eq.˜6.

### v2

@@ Abstract
Harnessing parallelism in seemingly sequential models is a central challenge for modern machine learning. Several approaches have been proposed for evaluating sequential processes in parallel using iterative fixed-point methods, like Newton, Picard, and Jacobi iterations. In this work, we show that these methods can be understood within a common framework based on linear dynamical systems (LDSs), where different iteration schemes arise naturally as approximate linearizations of a nonlinear recursion. Moreover, we theoretically analyze the rates of convergence of these methods, and we verify the predictions of this theory with several case studies. This unifying framework highlights shared principles behind these techniques and clarifies when particular fixed-point methods are most likely to be effective. By bridging diverse algorithms through the language of LDSs, the framework provides a clearer theoretical foundation for parallelizing sequential models and points toward new opportunities for efficient and scalable computation.
@@ 1 Introduction
Sequential processes are ubiquitous in machine learning models. Evaluating a recurrent neural network (Goodfellow et al., 2016), sampling a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), generating from a deep state space model (Gu et al., 2022; Smith et al., 2023; Orvieto et al., 2023; Gu and Dao, 2024), and unrolling layers of a deep neural network (He et al., 2016; Vaswani et al., 2017) all involve sequential computations. Naively, these sequential computations require time proportional to the sequence length or the network depth and do not take full advantage of hardware accelerators like GPUs and TPUs (Lim et al., 2024). However, for one important class of computations—linear recursions or linear dynamical systems (LDSs)—this bottleneck has been overcome using techniques like the parallel scan [^1] See Appendix B for an introduction to the parallel scan. (Blelloch, 1990; Fatahalian and Olukotun, 2024). Indeed, the parallelizability of linear recursions is key to scalably evaluating many deep state space models (Smith et al., 2023; Gu and Dao, 2024). A natural question is whether other sequential processes in machine learning, such as those with nonlinear recursions, can be similarly accelerated.
On first inspection, the parallel scan algorithm does not seem to generalize to nonlinear recursions. The parallel scan applies to linear recursions because the composition of two linear functions remains linear, whereas the composition of two nonlinear functions is generally more complicated. For example, the composition of two quadratic functions is quartic. Nevertheless, recent works have proposed several techniques to parallelize nonlinear recursions, including Jacobi (Song et al., 2021a), Picard (Shih et al., 2023; Lu et al., 2025), Newton (Danieli et al., 2023; Lim et al., 2024) and quasi-Newton (Tang et al., 2024; Gonzalez et al., 2024) iterations. These techniques were originally applied to different machine learning problems, including parallelizing the training of nonlinear RNNs and sampling from diffusion models, and follow different notations and intuitions, which obscures their underlying similarities.
In this paper, we show that Jacobi, Picard, and Newton iterations all parallelize nonlinear recursions over the sequence length by iteratively linearizing the nonlinear recurrence and evaluating the resulting linear dynamical system in parallel. While the connections between Picard and Newton iterations and their convergence rates are well-known in applied mathematics (Ortega and Rheinboldt, 1970), we make them explicit for nonlinear recursions and characterize when each method works and why.
@@
  • A unifying framework casting all four methods (Newton, quasi-Newton, Picard, and Jacobi) as iterative LDS evaluations (Section˜2);
  • A theoretical analysis showing that the rate of convergence depends on the stability of the resulting LDS and the fidelity of its approximation to the true linearized dynamics (Section˜3); and
  • Three case studies—the group word problem, evaluating a nonlinear RNN, and sampling from a discretized Langevin diffusion—validating that this analysis predicts which method suits which problem (Section˜4).
These contributions unify and clarify several recently proposed methods in the machine learning literature, and they highlight the centrality of linear dynamical systems for parallelizing seemingly sequential processes.
@@ 2 Unifying fixed-point iterations using linear dynamical systems
We first introduce notation for evaluating a generic sequence model. Let $x_{t}\in\mathbb{R}^{D}$ denote the state at time $t$, and let $f_{t}$ denote the corresponding transition function at that time point. Throughout this paper, we will use $D$ to denote the dimension of the hidden state and $T$ to denote the sequence length.
@@
In order to be a fixed-point operator, $\mathcal{A}$ should have a fixed point $\mathbf{x}_{1:T}^{\star}$. That is, $\mathcal{A}$ should satisfy that $\mathbf{x}_{1:T}^{\star}=\mathcal{A}\left(\mathbf{x}_{1:T}^{\star}\right)$, where $\mathbf{x}_{1:T}^{\star}$ is the unique root of the system of equations given by
$x_{t+1}-f_{t+1}(x_{t})=0,\quad\forall t\in\{0,\ldots,T-1\}.$ (3)
@@
$x_{t+1}^{(i+1)}$ $=f_{t+1}(x_{t}^{(i)})+\tilde{A}_{t+1}\,(x_{t}^{(i+1)}-x_{t}^{(i)})$ (4)
where the transition matrix $\tilde{A}_{t+1}\in\mathbb{R}^{D\times D}$ is determined by the dynamics functions $f_{t+1}$ and the current guess for the state $x_{t}^{(i)}$. Importantly, eq.˜4 is a linear dynamical system for the as yet unknown $\textbf{x}_{1:T}^{(i+1)}$ in terms of the currently known $\textbf{x}_{1:T}^{(i)}$, as $x_{t+1}^{(i+1)}$ is a linear function of $x_{t}^{(i+1)}$ plus a bias $b_{t+1}=f_{t+1}(x_{t}^{(i)})-\tilde{A}_{t+1}x_{t}^{(i)}$, which only depends on the states from the previous fixed-point iteration.
@@
Table 1: Summary of fixed-point iteration schemes as linear dynamical systems. We list the methods by the order of their approximation. While higher order methods may converge in fewer iterations, each iteration may be more costly. For example, the prefix sum and parallel scan have $\mathcal{O}(\log T)$ depth, while a single Jacobi iteration has constant depth. For all the methods, each iteration is an LDS, i.e. they can be written in the form of eq.˜4 where $\tilde{A}_{t+1}$ is the transition matrix. These methods are guaranteed to converge in at most $T$ iterations (Gonzalez et al., 2024, Proposition 1). “Order” means the highest number of derivatives taken: Newton and quasi-Newton methods use first derivatives, while Picard and Jacobi methods do not use derivatives of $f_{t}$.
@@
Fixed-point method Order Transition matrix $\bm{\tilde{A}_{t+1}}$ Parallelization
Newton first-order $\dfrac{\partial f_{t+1}}{\partial x_{t}}(x_{t}^{(i)})$ Parallel Scan (dense matrix multiplication)
@@
Quasi-Newton quasi first-order $\operatorname{diag}\!\left[\dfrac{\partial f_{t+1}}{\partial x_{t}}(x_{t}^{(i)})\right]$ Parallel Scan (elementwise vector multiplication)
Picard zeroth-order $I_{D}$ Prefix Sum (vector addition)
Jacobi zeroth-order 0 Map (embarrassingly parallel)
Figure 1: Parallel scan for LDS. We illustrate the parallel scan for a sequence of length $T=4$ for the simple LDS $x_{t}=A_{t}x_{t-1}$.
The transition matrix $\tilde{A}_{t+1}$ can be thought of as an approximation to the Jacobian of the dynamics function $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. Different fixed-point methods simply use different approaches to linearizing the dynamics function, as shown in Table˜1.
Importantly, because the recursion in eq.˜4 is an LDS, it can be evaluated with a parallel scan (Blelloch, 1990; Nguyen, 2007). The parallel scan is a core primitive that allows the unrolling of $T$ steps from an LDS, $x_{t+1}=A_{t+1}x_{t}+b_{t+1}$, in $\mathcal{O}(\log T)$ time on a machine with $\mathcal{O}(T)$ processors. The logarithmic time in the sequence length comes from a divide-and-conquer approach made possible from the fact that composition of affine functions is closed, i.e. another affine function. To show the closure of composition of affine functions more explicitly, note that if $f_{t}(x)=A_{t}x+b_{t}$ and $f_{t+1}(x)=A_{t+1}x+b_{t+1}$, then $f_{t+1}(f_{t}(x))=A_{t+1}A_{t}x+\left(A_{t+1}b_{t}+b_{t+1}\right)$. Therefore, as shown in Figure˜1, by combining neighboring affine functions together, we obtain transition maps from the initial condition $x_{0}$ to all other states $x_{t}$ in $\mathcal{O}(\log T)$ time. We provide further discussion of the parallel scan algorithm for the interested reader in Appendix˜B.
@@
Next, we discuss in more detail how the four prominent fixed-point methods discussed in Table˜1 reduce to iterative application of LDSs when used to solve a recursion. Fundamentally, all of these methods parallelize nonlinear recursions by iteratively linearizing and evaluating them, as we indicate in Algorithm˜1. However, the different methods use different approximations of the Jacobian $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ of the dynamics function for their transition matrices $\tilde{A}_{t+1}$.
@@
   $\tilde{A}_{1:T}\leftarrow\textsc{LinearizeDynamics}(f,x_{0},\mathbf{x}_{1:T}^{(i)})$ $\triangleright$ For all $t$ in parallel
   $\mathbf{x}_{1:T}^{(i+1)}\leftarrow\textsc{EvaluateLDS}(x_{0},\tilde{A}_{1:T},f,\mathbf{x}_{1:T}^{(i)})$ $\triangleright$ Using parallel scan
@@ 2.1 Newton iterations
Bellen and Zennaro (1989), Gander and Vandewalle (2007), Danieli et al. (2023), and Lim et al. (2024) demonstrated that when the fixed-point operator $\mathcal{A}$ is constructed as an appropriately designed LDS, built from a linearization of original nonlinear recursion $f$, then each application of $\mathcal{A}$ is equivalent to an iteration of Newton’s root-finding method on the system of equations given in eq.˜3. Specifically, each Newton fixed-point iteration, $\mathbf{x}_{1:T}^{(i+1)}=\mathcal{A}_{\mathrm{N}}\big(\mathbf{x}_{1:T}^{(i)}\big)$}, is defined by the linear recursion,
@@
which we recognize as the first-order Taylor expansion of the nonlinear recursion. Because this fixed-point iteration uses a first derivative, we refer to it as a first-order fixed-point iteration. We note that for Newton iterations, the transition matrix is exactly the Jacobian of the dynamics function, $\tilde{A}_{t+1}:=\tfrac{\partial f_{t+1}}{\partial x_{t}}(x_{t}^{(i)})$.
Crucially, because eq.˜5 is a linear dynamical system, it can be evaluated in parallel over the sequence length using an associative scan (Blelloch, 1990). However, each iteration requires $\mathcal{O}(TD^{2})$ memory to store the $T$ Jacobian matrices, and $\mathcal{O}(TD^{3})$ work to compute the matrix-matrix multiplications in the parallel scan. This expense is prohibitive for large state size or sequence length, which motivates the quasi-Newton iterations we discuss next.
@@ 2.2 Quasi-Newton iterations
The compute and memory costs of full Newton iterations have motivated a wide literature on quasi-Newton methods (Nocedal and Wright, 2006). A particularly simple way to turn the Newton iteration in eq.˜5 into a quasi-Newton iteration that uses a parallel associative scan was proposed by Gonzalez et al. (2024): just use the diagonal of the Jacobian of the dynamics function. Specifically, each of these quasi-Newton fixed-point iterations $\mathcal{A}_{\mathrm{QN}}$ is given by
@@
which we recognize as an LDS with transition matrix $\tilde{A}_{t+1}=\mathrm{diag}\big[\tfrac{\partial f_{t+1}}{\partial x}(x_{t}^{(i)})\big]$.
With this transition matrix, the parallel scan requires only $\mathcal{O}(TD)$ memory and $\mathcal{O}(TD)$ work, and is therefore more computationally efficient than the full Newton iteration eq.˜5. However, quasi-Newton methods may take more fixed-point iterations to converge. As we will show in Sections 3 and 4, whether Newton and quasi-Newton methods are faster for evaluating a given sequence depends on many factors including the accuracy of the approximate Jacobian, the stability of the linearized system, the choice of hardware, and the scale of the problem. In many cases, the large memory consumption of full Newton iterations renders it infeasible for large-scale problems.
Another consideration is the manner in which $\tilde{A}_{t+1}$ is computed. A primary goal of quasi-Newton methods (Broyden, 1970; Dennis Jr and Schnabel, 1996) is to avoid computation of $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. Simply computing these Jacobians with autodifferentiation and then taking their diagonal still requires $D$ function calls, which can be costly. However, in sequence modeling with recurrent neural networks (RNNs), the diagonal elements of the Jacobian can often be written in closed form, allowing for its evaluation in a single function call. This approach is taken in Gonzalez et al. (2024) and Danieli et al. (2026). Another approach proposed in Zoltowski et al. (2025) is to stochastically estimate the diagonal using the Hutchinson estimator. This stochastic approach also requires only one function call but is generally applicable even when closed forms of $\tilde{A}_{t+1}$ are not available. We employ this efficient stochastic estimator in all of our experiments. Finally, a third approach comes from the multisecant or Broyden family of methods (Fang and Saad, 2009; Walker and Ni, 2011), including Anderson acceleration (Anderson, 1965; Bai et al., 2019; Tang et al., 2024), which build up estimates of derivative information over the fixed-point iterations.
In general, any structured approximation of the Jacobian matrix that remains closed under matrix multiplication [^2] Closed in the sense that the product, $A_{t+1}A_{t}$, has the same structure as $A_{t+1}$ and $A_{t}$, as with diagonal matrices. See Appendix B for more detail. could be used to form a quasi-Newton fixed-point iteration amenable to a parallel scan. For example, Zoltowski et al. (2025) introduces a parallel quasi-Newton method where each block of the matrix is diagonal. The broader development of quasi-Newton methods that fit into the unifying LDS framework discussed in this paper is an important direction for future work, which we elaborate on in Section˜6. However, for simplicity, henceforth in this paper when we refer to “quasi-Newton iterations,” we restrict ourselves to the simple diagonal approximation shown in eq.˜6.